OpenStudy (anonymous):

\[\lim_{x \rightarrow 0} \frac{ e^{x^{2}}-1 }{ 2x^{2} }\] find the limit as x aproaches zero

OpenStudy (anonymous):

are you familiar with l'Hopital's Rule?

OpenStudy (zzr0ck3r):

do you know l'hopitals rule?

OpenStudy (anonymous):

yes 0/0 or 1/1 right?

OpenStudy (zzr0ck3r):

do it twice

OpenStudy (anonymous):

0/0 or inf/inf. 1/1 is fine, limit exists and = 1

OpenStudy (anonymous):

2-1/4 = 1/4

OpenStudy (zzr0ck3r):

?

OpenStudy (anonymous):

;( that's not it

OpenStudy (zzr0ck3r):

you will get (4x^2*e^x^2)/4 --> 0

OpenStudy (anonymous):

oh ty

OpenStudy (anonymous):

i get limit = 1/2

OpenStudy (anonymous):

1/2, or zero

OpenStudy (zzr0ck3r):

second derivative of the top is 4x^2 * e^x^2 right?

OpenStudy (zzr0ck3r):

second on bottom is 4

OpenStudy (zzr0ck3r):

o im out of it, I didn't use product rule...

OpenStudy (zzr0ck3r):

ignore me please

OpenStudy (zzr0ck3r):

I don't have paper and im lazy:)

OpenStudy (anonymous):

first time \[\frac{ 2xe^{x^{2}} }{ 4x } \] second time \[\frac{ 2e^{x^{2}} + 4x^{2}e^{x^{2}} }{ 4 }\]

OpenStudy (anonymous):

yup. only thing was product rule

OpenStudy (anonymous):

can i start another thread so i can give you and metal too Euler. Because you both are super helpfull :)